OpenStudy (anonymous):

Find the exact value of cos(19pi/12)

OpenStudy (jdoe0001):

well, notice that 19/12 is really 10/12 + 9/12 which simplified will be 5/6 + 3/4 so use that :)

OpenStudy (jdoe0001):

and you'd surely have those in your Unit Circle

OpenStudy (anonymous):

that would make it 285 degrees

OpenStudy (jdoe0001):

well, it'll be a value, the cosine, no the arcCosine, the angle itself

OpenStudy (anonymous):



OpenStudy (jdoe0001):

\(\bf cos\left(\cfrac{19\pi}{12}\right) \implies cos\left(\cfrac{10\pi}{12}+\cfrac{9\pi}{12}\right) \implies cos\left(\cfrac{5\pi}{6}+\cfrac{3\pi}{4}\right)\\ cos(a+b) = cos(a)cos(b) - sin(a)sin(b)\)

OpenStudy (anonymous):

cos(5pi/6) cos(3pi/4) - sin(5pi/6) sin(3pi/4)

OpenStudy (anonymous):

How do I go from there?

OpenStudy (jdoe0001):

those are common values, I assume you have a Unit Circle?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

then just get the values off the Unit Circle :)

OpenStudy (anonymous):

do you mean the degrees?

OpenStudy (jdoe0001):

the cosine and sine values for the angle, maybe you need a better unit circle, with radians and cosine and sine values => http://i.stack.imgur.com/r8uHr.gif

OpenStudy (jdoe0001):

as you can see in the upper-right-hand corner, the value pair is the (cos, sin)

OpenStudy (anonymous):

I'm still confused

OpenStudy (jdoe0001):

ok.... what confuses you?

OpenStudy (anonymous):

How do you get the value of cos(5pi/6) from the unit circle?

OpenStudy (anonymous):

actually I get it

OpenStudy (anonymous):

[-squareroot of (3)/2 times -squareroot of (2)/2] - [1/2 times squareroot of (2)/2]

OpenStudy (jdoe0001):

5pi/6 is really just 150 degrees in the unit circle next to that, you'll see a pair like => \( \bf \left(-\cfrac{\sqrt{3}}{2},\cfrac{1}{2}\right)\)

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

square root of (6)/4 - square root of (2)/4

OpenStudy (jdoe0001):

\(\left(-\cfrac{\sqrt{3}}{2},\cfrac{1}{2}\right) \implies (cos,sin)\)

OpenStudy (anonymous):

\[[\sqrt{6}-\sqrt{2}]/4\]

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thanks for helping

OpenStudy (jdoe0001):

yw